OpenStudy (buggiethebug):

What is the maximum number of real distinct roots that a quartic equation can have?

OpenStudy (anonymous):

An nth degree polynomial can have as many as n real roots.

OpenStudy (anonymous):

The number of distinct real roots 1 of the cubic equation x 3 + bx 2 + cx + d = 0 (1) equals, for example, the number of normals to the parabola y = x 2 through a given point in the plane (Bains and Thoo (2007)), 2 as well as the number of equilibrium solutions of dx/dt = x 3 + bx 2 + cx + d . Now, one way to find the number of real roots of (1) is to solve th e equation. Certainly, (1) can be solved by hand easily if d = 0. If d 6 = 0, the equation can still be solved by hand using Cardan’s (or Cardano’s) cubic f ormula (Fine (1961); Gellert et al. (1975)), 3 but not easily in general. Of course, using computer software or a graphing calculator can make light work of solv ing (1) altogether. However, there is a certain satisfaction in being able to tel l the number of real roots of (1) without first solving the equation. It turns out t hat this is easy to do. The key is the discriminant . Every intermediate algebra student learns the quadratic fo rmula, deriving it by completing the square. And with the quadratic formula i n hand, it is apparent that the number of real roots of the quadratic equat ion x 2 + bx + c = 0 ∗ 1501 Mendocino Avenue, Santa Rosa, CA 95401-4332, rkavinoky@santarosa.edu . † 2088 N. Beale Road, Marysville, CA 95901-7605, jthoo@yccd.edu . 1 We mean the number of distinct real roots throughout. 2 It is remarkable that Apollonius had obtained precise resul ts on the number of normals to a parabola through a given point using purely synthetic geom etry (Heath, 1981, pp. 158–159, 163–166) almost 1700 years before Italian mathematicians s olved the cubic. 3 For Cardan’s formulation of the solution of the cubic, as wel l as the story behind the quarrel between Cardan and Tartaglia over its publication, see Burton (2003) for example. For other formulations of the solution of the cubic, see Kalm an and White (1998) and the references therein.